The area of ABCD is 90 square inches. The length is ten times more then its width. What is the length? What is the width? What is the length of the diagonal?
Answer by Alan3354(69443)

The area of ABCD is 90 square inches. The length is ten times more then its width. What is the length? What is the width? What is the length of the diagonal?
--------------
Area = L*W = 90
L = 11W (10 times more = 11 times)
Sub for L in the 1st eqn
11W*W = 90
W^2 = 90/11
Width = sqrt(90/11) =~ 2.86 inches
Length =~ 31.464 inches
------------------------